PartnerObserving Log for 2007 June 21-22

Observer: J. Hill
LBC Support: A. DiPaola, R. Speziali, R. Bertram
Telescope Operator: D. Gonzalez

More technical observing with SDT OBs and assorted other projects mixed in.

Clouds and even some light rain showers in the late afternoon. Clearing after sunset. Tried to open at 03:30 UT, but the PLC software is having some issues with the shutter doors (3 of the locked out latch pins are in an undefined state). See Issue #700. Solved our PLC issues and opened at 05:24 UT. T=14degC. D=-2degC Wind 3m/sec from North. Scattered clouds.

5 bias

Spent half an hour resetting elevation encoders.

We'll try some SDT OBs just to be sure that everything is put back together correctly. While I am observing, Andrea is logging various error/debugging messages.

06:00 - Vfastextra at position of M92, but just now we are looking through clouds so no pupils were found.

06:15 - Vfastextra 32 sec: 2068 0 0 0 0 0 sent

06:18 - Vfastextra 32 sec: 124 -392 -324 -1614 948 -21 (not sent, used manual +300 Z11)

06:21 - Vfastextra 32 sec: -649 138 -510 -429 -145 358 sent

06:23 - Vfastextra 32 sec: -332 -481 175 -396 194 346 sent

06:26 - Vfastextra 32 sec: -543 -318 -293 5 -122 18 sent

06:28 Ferraro M92_deepR_F2_6 - Guidestar for dither 1 is elliptical left-right implying that we have some astigmatism and are not well focussed. It improves as the exposure progresses. Images are around 1 arcsec and showing the astigmatism. Sky background is 12000 counts in 300 sec (moon still up). Images improved to 0.7 arcsec in the second dither. There are broken clounds drifiting by, so nothing is photometric about these images.

06:48 Vfastextra: -792 317 206 19 -43 -145 sent

06:50 M92_deepR_F2_7 - Guidestar is still astigmatic. Adusting the focus manually during dither 1. Images are around 0.7 arcsec limited by alignment/focus rather than by seeing.

Move to field of M16

07:14 Vfastextra: -1270 689 168 -183 558 54 sent

07:16 Vfastextra: -376 275 103 21 140 101 sent

Pupils are beautiful here, but we have reached the focus limit in +Z because of the warm temperature. I guess I'll have to move the Blue hub tomorrow.

07:19 m16_1Ugr - Running Dave's OB slightly defocussed as it will be useful to gage the exposure levels. Beware that we might be looking through some thin clouds or smoke this low in the south. r' image is remarkably pretty, although still slightly out of focus. Yes, the red flag on reaching the Z-limit does work.

Move to field of GRB070518

07:33 Vfastextra: -218 -1144 -139 185 -419 -227 sent

07:35 Vfastextra: 73 -192 203 140 -66 -156 sent

07:39 grb070518 - 1 arcsec images are limited by aberrations in dither 1. Seeing exploded to 3 arcsec during dither 3. Sky is mostly clear. Moon is down. Dither 4 has 1.4 arcsec images limited by seeing. Discovered that PCS guiding was turned off during dither 5. This explains some og the elliptical science images, but not the elliptical guide stars we saw earlier.

08:00 grb070518 second pass - guide star still has fluffy look (2 arcsec) in dither 1, but slowly improving.

Move to field of SS433

08:28 Vfastextra: 415 625 249 85 179 -44 sent

08:31 Vfastextra: -31 9 -311 52 -6 -55 sent

08:35 SS433e - Images are back to 0.7 arcsec or better in r at 1.13 airmasses. We see astigmatism at the bottom of chip 2, so it seems like my field astigmatism corrections from last week are not working as well as they did. The correct numbers seem to be in PSF. This field has an ungodly number of stars in it! We see plenty of nebular emission in Chip 3, so will run the second OB.

09:10 Vfastextra: 110 32 -249 105 55 10 sent

09:13 SS433_pa90 - Images in r' approaching 0.6 arcsec at 1.15 airmasses. Science.py is crashing DS9 every couple of frames. Focus is better now, so astigmatism is barely visible.

Move back to field of GRB070518

09:50 Vfastextra: -201 146 174 30 77 5 sent (note small change from previous pointing on the other side of the sky)

09:53 GRB070518 third pass - Saw image jump by an arcsec near 02:54:16 MST. Dither 1 is 0.8 arsec in r' at 1.40 airmasses. Image jumps again near 03:04:00 MST.

10:xx GRB070518 fourth pass - Now I'm fiddling with the guiding more than doing science.

10:21 Changed guide .dat to the following during second dither:
TypicalExposureTime=4 (was 8)
MaximumExposureTime=32
MinGuidingCorrection=0.01
MaxTrackingErrorPerSec=0.1 (was 0.25)
MinimumTrackingStars=1
AzTrackingGain=0.8 (was 1.0)
ElTrackingGain=0.8 (was 1.0)

10:45 STOP OB after it has been frozen on dither 5 for more than 10 minutes. Note that UI was not frozen - at least the clock was counting. Is this related to the LBC archive being frozen? OB frees at 10:52 while I was fiddling with space in /var on CMU.

TCS is doing strange things about the same time. Preset and MCS seem to be frozen.

11:00 Start taking SkyFlatG_5 blind, but the UI says "stop" at the bottom. ?????

11:03 Stop and restart OB. But apparently PCS was hung too. MCSPU is not getting any polynomials.

11:15 Rebooted lbcdh1=lbcarchive2 (Riccardo could not connect to it)

11:18 Found the stray LSS that was blocking PCS. Probably LSS was started when David pressed the PMCR launch button. This (PCS) in turn unblocked the OB.

11:20 Riccardo and / or the reboot gets the archive running again. Surely that wasn't blocked by LSS?

11:24 SkyFlatG_rot90 x1 - Dither 2 has the WCS error in header. Stopped because exposures too short.

11:27 SkyFlatG_rot90 x6

11:33 SkyFlatG_rot180 x3

11:38 SkyFlatG_rot270 x1 (dither 5 saturated)

Sky is not photometric, but was clear in the direction of the flats. These were special 4-rotation flats for Vincenzo.

Close Dome T=12.6, D=-3.2 Wind 1 m/sec from N.

11:48 25 bias

-- JohnHill - 22 Jun 2007
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback